# Thermodynamic properties of the macroscopically degenerate tetramer-dimer phase of the spin-1/2 Heisenberg model on the diamond-decorated square lattice

### v1

@@ III.4 Quantum Monte Carlo
More specifically, our QMC approach is based on the stochastic series expansion (SSE) method with directed loop updates [66, 67, 68]. In order to avoid the sign problem, i.e., an exponential drop of the statistical accuracy of the QMC simulations at low temperatures and large system sizes [29, 30, 69, 70] due to the presence of geometric frustration, we avoid working in the conventional local spin-$S^{z}$ basis for the model at hand. Instead, we eliminate the sign problem upon using appropriate basis states after decomposing the Hamiltonian into separate terms based on dimers (or trimers) [71, 72, 73, 74]. For certain so-called fully frustrated models the sign problem can indeed be completely eliminated, cf. Refs. [72, 73, 75, 76, 77] for the spin-dimer and Ref. [74] for the spin-trimer basis, respectively.
In case of the diamond-decorated square lattice a finite value of the coupling $J_{2}$ leads to geometric frustration. In order to avoid the associated sign problem, we follow Ref. [48] and treat all $J_{2}$-dimer spins in the spin-dimer basis, but keep using the local $S^{z}$-basis for the monomer spins. Using this combined five-site cluster-basis, we can simulate our model without a sign-problem within the SSE framework, based on the abstract operator loop update introduced in Ref. [74]. We refer to Ref. [74] for further details on this QMC approach. Here, we performed QMC simulations for systems with up to $6\times 6$ unit cells.
@@ IV.1 Dimer triplet density, entropy, and specific heat above the DT phase
Figure 8 presents results for the interaction ratio $J_{2}/J_{1}=1.7$. At zero temperature, the dimer triplet density is 0.25, reflecting the nature of the DT state, where $1/4$ of the dimers are in the dimer-triplet state within the singlet tetramers and $3/4$ of the dimers remain in the dimer-singlet state. Due to the relatively lower energy cost of creating a dimer-singlet state compared to an additional dimer-triplet state for the interaction ratio $J_{2}/J_{1}=1.7$, a decrease in the dimer-triplet density can be observed at low temperatures terminated at a round local minimum. With further increase of temperature, the dimer-triplet density tends towards the asymptotic value $0.75$ reached in the limit of infinite temperature due to the three-fold degeneracy of the triplet state in zero magnetic field.
A comparison of the dimer-triplet density obtained for $N_{s}=30$ and $N_{s}=40$ spins from exact diagonalization (ED) reveals small finite-size effects. ED for $N_{s}=40$ spins provides an interpolation between the spin-star decoupling approximation and the EMD model at low temperatures and QMC at high temperatures. The excitation from a dimer-singlet state to a dimer-triplet state is neglected within the EMD model and the dimer-triplet density should accordingly always decrease with increasing temperature. On the other hand, the local minimum observed in the dimer-triplet density arises from the competition of the lower excitation energy for the dimer-singlet state and the higher degeneracy of the triplet state. In fact, the lowest states not taken into account by the EMD model are at energies slightly above $0.6J_{1}$ for $J_{2}=1.7J_{1}$ and these are responsible for the triplet density increasing again with increasing temperature after going through a minimum. However, the spin-star decoupling approximation takes only some of these into account (as reflected also by the lower entropy $s$ compared to ED and QMC, see Fig. 8(b)) and approximates their energies, thus explaining the deviation from $N_{s}=40$ ED and QMC, which are in good agreement for $T/J_{1}\gtrsim 0.2$.
@@
Figure 8(c) displays the specific heat per spin $c$ as a function of temperature for the same value of the interaction ratio. A notable feature is the presence of a low-temperature peak respectively shoulder superimposed on the high-temperature maximum of the specific heat. As anticipated, the spin-star and EMD approaches fail to capture the behavior of the specific heat at higher temperatures as they neglect high-energy excitations above the DT phase. On the other hand, these approaches accurately account for low-lying excitations which leads to a close agreement between the ED data for $N_{s}=40$ and the analytical results obtained from the spin-star decoupling approximation at low enough temperatures. A comparison of the peak height in the specific heat obtained from ED and QMC reveals relatively large finite-size effects at low temperatures. The EMD model provides further insight into the behavior of the specific heat at low temperatures, which effectively bridges the gap between the available QMC data ($T/J_{1}>0.1$) and extends it further towards lower temperatures. Figure 8(c) shows that thermal activation of the specific heat indeed occurs at progressively lower temperatures as the system size increases such that the EMD model provides the best approximation to the thermodynamic limit at low temperatures.
@@
We now turn to a discussion of the dimer-triplet density, entropy, and specific heat in zero magnetic field for the case $J_{2}/J_{1}=1.5$ plotted in Fig. 9. Here the numerical data for the dimer-triplet density exhibits only a shallow minimum, which is consistent with the gap for one triplet less being only slightly smaller than that for an additional triplet at $J_{2}/J_{1}=1.5$, compare Fig. 7. By contrast, the spin-star decoupling approximation fails to account for this minimum. The reason is that, within the spin-star decoupling approximation, the excitation energies from the DT phase to singlets and triplets are equal for this interaction ratio. Consequently, the spin-star approach cannot explain the thermal population into singlet states for the interaction ratio $J_{2}/J_{1}\leq 1.5$. Comparing the results obtained from QMC simulations for a system size of $6\times 6$ (total of 180 spins) with those from ED results for 40 spins reveals nearly perfect agreement, indicating negligible finite-size effects. However, the discrepancy between ED results for 30 spins and 40 spins at intermediate temperatures ($T/J_{1}>0.1$) may stem from artificial symmetries imposed for the system size of 30 spins by considering periodic boundary conditions, which are absent for the system size of 40 spins.
The entropy and specific heat for the interaction ratio $J_{2}/J_{1}=1.5$ in zero magnetic field are shown in Fig. 9(b) and 9(c), respectively, and exhibit qualitatively similar behavior to that for the interaction ratio $J_{2}/J_{1}=1.7$. In the high-temperature regime, small finite-size effects are observed. While the high-temperature peak of the specific heat shifts slightly towards higher values with increasing system size, the low-temperature peak is conversely suppressed towards lower values. The results obtained from the spin-star decoupling approximation for 40 spins are slightly overestimated compared to ED data for $N_{s}=40$ at low temperatures $T/J_{1}\lesssim 0.1$. The discrepancy at higher temperatures arises due to the presence of low-energy excitations to states that share the same energy as the Lieb-Mattis ferrimagnetic state at zero magnetic field (compare the discussion in Sec. III.3), which is not accurately accounted for by the spin-star decoupling approximation simply ignoring the non-commutativity of the spin-star Hamiltonians. The QMC data are consistent with the intriguing double-peak temperature dependence of the zero-field specific heat with low- and high-temperature peaks located around $T/J_{1}\approx 0.12$ and 0.6, respectively.
@@
With a yet lower value of the interaction ratio $J_{2}/J_{1}$, the analytical approaches provided by the spin-star and EMD models neglect more and more important low-energy excitations. For the interaction ratio $J_{2}/J_{1}$=1.3, both of these approaches only provide an accurate description at very low temperatures as evidenced by the temperature dependencies of the dimer-triplet density, entropy, and specific heat depicted in Fig. 10. This limitation arises from the failure of these analytical approaches to accurately capture the degeneracy of the LM phase in zero magnetic field with possible values of the $z$-component of the total spin $S_{T}^{z}=-S_{T},-S_{T}+1,...,S_{T}$, and the low-energy excitations to these states are not properly accounted for either. The dimer-triplet density monotonically increases without any local minimum for the interaction ratio $J_{2}/J_{1}$=1.3, as evidenced by the ED data for $N_{s}=40$ spins. This behavior is due to the lowest excitation in this parameter regime arising from an additional dimer-triplet, compare Fig. 7. The EMD approach neglects the lowest excitations and thus only gets the $T=0$ limit right, but fails to capture the leading low-temperature asymptotics. By contrast, the star approach does account for these excitations, even if only in an approximate fashion, and thus captures at least qualitatively the low-temperature behavior of the triplet density.
@@ IV.2 Magnetization, entropy, and specific heat under magnetic field
It can be observed from Fig. 11(b) that the entropy starts from its nonzero residual value, which is progressively converging with increasing system size to the residual entropy of the dimer model on the square lattice per unit cell $s_{d}\approx 0.29156$ (per spin $\approx 0.058312$) [64, 65, 43]. Contrary to this, one detects a sizable entropy gain reaching the specific value $s=0.13256$ at the transition field between the DT and MD phases, which is almost independent of the system size as corroborated by ED, the EMD model, as well as the spin-star decoupling approximation. Similar qualitative agreement between results of all three aforementioned methods is observed in the peculiar asymmetric double-peak structure of the specific heat, which can be detected in the vicinity of the magnetic-field-driven phase transition between DT and MD phases (see the inset of Fig. 11(c)). The lower peak height observed below the relevant transition field can be attributed to the macroscopic degeneracy of the DT ground state, which is not lifted by the magnetic field in contrast to the MD phase that becomes non-degenerate at any nonzero magnetic field.
@@
The results obtained from the spin-star decoupling approximation are in very good agreement with the ED data even for magnetic fields higher than $h/J_{1}\gtrsim 1.0$, up to the end of the Lieb-Mattis ferrimagnetic phase, and even at moderately high temperatures $T/J_{1}=0.1$ still constitute a good approximation of the field dependence of the entropy and specific heat, as depicted in Fig. 11(b) and (c).
@@
Figure 12: (a) Magnetization; (b) entropy; (c) specific heat for the interaction ratio $J_{2}/J_{1}=1.5$ as a function of magnetic field at two different values of temperature $T/J_{1}=0.01$ and $T/J_{1}=0.1$. The inset of panel (a) shows $T=0$ ED and DMRG results.
@@
Figure 12(a) shows the magnetization curves for the interaction ratio $J_{2}/J_{1}=1.5$. There is a distinct magnetization jump from the DT phase to the MD phase at zero temperature, as shown in the inset of Fig. 12(a). However, the 1/5-plateau is sufficiently small to be smeared out at finite temperatures. Consequently, there is a sharp rise in magnetization from the 0-plateau to the 3/5-plateau even at temperatures as low as $T/J_{1}=0.01$, which becomes smoother with increasing temperature. Due to the presence of the MD phase within a narrow range of magnetic-field strengths, the EMD model is in excellent agreement with the ED data up to a magnetic field of approximately $h/J_{1}\approx 0.54$ at a very low temperature $T/J_{1}=0.01$ though it fails to account for further increases in magnetization with increasing magnetic-field strength. On the other hand, the spin-star decoupling approximation is still capable of qualitatively reproducing the ED data even at higher magnetic fields. The comparison of ED data for $N_{s}=40$ spins with QMC data for 36 unit cells (i.e., $N_{s}=180$ spins) indicates negligible finite-size effect in this parameter region.
@@
The results for the low-temperature entropy exhibit a relevant discrepancy between the ED and spin-star decoupling approximation only in the close vicinity of the transition field $h/J_{1}=0.5$ (see Fig. 12(b)). The excessive macroscopic degeneracy predicted by the spin-star decoupling approximation relates to a triple phase coexistence point of the LM, DT, and MD phases, which is located exactly at the magnetic field $h/J_{1}=0.5$ for $J_{2}/J_{1}=1.5$. Recall that the triple point obtained within the spin-star approach is slightly shifted towards a lower value of the magnetic field with respect to its true coordinates (see Fig. 2(b) and the discussion in Sec. III.3). At a higher temperature $T/J_{1}=0.1$ the discrepancy between ED and the spin-star decoupling approximation is significantly reduced, because all relevant low-lying excited states are thermally activated. Contrary to this, we see the opposite effect for the EMD model with much lower entropy, because high-energy excitations are missing within this description. A similar effect can also be observed in the field dependence of the specific heat (Fig. 12(c)). While the EMD model reproduces the ED data at low temperatures, the spin-star decoupling approximation correctly describes the double peak of the specific heat also at a higher temperature of $T/J_{1}=0.1$. At this higher temperature, the double-peak structure of the specific heat is also confirmed by the QMC simulations for the bigger system of $6\times 6$ unit cells.
@@
Last but not least, magnetization, entropy, and specific heat as a function of temperature are depicted in Fig. 13 for the interaction ratio $J_{2}/J_{1}=1.3$. As discussed in Sec. III.3, for the interaction ratio $J_{2}/J_{1}\lesssim 1.4$, the lowest excitation above the DT phase corresponds to the creation of an additional triplet (see Fig. 7). For this reason, the EMD model is unable to describe the finite-temperature properties for the interaction ratio $J_{2}/J_{1}=1.3$. On the other hand, our second analytical approach derived from the spin-star decoupling approximation provides a quite reliable description of all quantities above the transition field $h/J_{1}=0.3$. The most important artifacts of the spin-star decoupling approximation are just below the transition field and at $T/J_{1}=0.01$ where it yields a spike in the entropy $s$ (Fig. 13(b)) and a second maximum in the specific heat $c$ (Fig. 13(c)) while such pronounced maxima are absent in the ED results. The reason for this difference is that the field-induced DT-LM transition is a true first-order transition in the full model such that excitations are scattered over a certain energy range below the transition while the spin-star decoupling approximation treats these as a degenerate manifold that comes down at the transition field. Conversely, a temperature $T/J_{1}=0.1$ is sufficiently high to thermally excite these low-lying excitations in the full model such that at this higher temperature the $N_{s}=40$ ED results also exhibit a pronounced maximum below the transition field, both in the entropy (Fig. 13(b)) and the specific heat (Fig. 13(c)). The latter behavior is confirmed by our QMC simulations on the bigger system of $6\times 6$ unit cells.

### v2

@@ III.4 Quantum Monte Carlo
More specifically, our QMC approach is based on the stochastic series expansion (SSE) method with directed loop updates [66, 67, 68]. In order to avoid the sign problem, i.e., an exponential drop of the statistical accuracy of the QMC simulations at low temperatures and large system sizes [29, 30, 69, 70] because of the presence of geometric frustration, we avoid working in the conventional local spin-$S^{z}$ basis for the model at hand. Instead, we eliminate the sign problem upon using appropriate basis states after decomposing the Hamiltonian into separate terms based on dimers (or trimers) [71, 72, 73, 74]. For certain so-called fully frustrated models the sign problem can indeed be completely eliminated, cf. Refs. [72, 73, 75, 76, 77] for the spin-dimer and Ref. [74] for the spin-trimer basis, respectively.
In case of the diamond-decorated square lattice a finite value of the coupling $J_{2}$ leads to geometric frustration. In order to avoid the associated sign problem, we follow Ref. [48] and treat all $J_{2}$-dimer spins in the spin-dimer basis, but keep using the local $S^{z}$ basis for the monomer spins. Using this combined five-site cluster basis, we can simulate our model without a sign problem within the SSE framework, based on the abstract operator loop update introduced in Ref. [74]. We refer to Ref. [74] for further details on this QMC approach. Here, we performed QMC simulations for systems with up to $6\times 6$ unit cells.
@@ IV.1 Dimer triplet density, entropy, and specific heat above the DT phase
Figure 8 presents results for the interaction ratio $J_{2}/J_{1}=1.7$. At zero temperature, the dimer triplet density is 0.25, reflecting the nature of the DT state, where $1/4$ of the dimers are in the dimer-triplet state within the singlet tetramers and $3/4$ of the dimers remain in the dimer-singlet state. Because of the relatively lower energy cost of creating a dimer-singlet state compared to an additional dimer-triplet state for the interaction ratio $J_{2}/J_{1}=1.7$, a decrease in the dimer-triplet density can be observed at low temperatures terminated at a round local minimum. With further increase of temperature, the dimer-triplet density tends towards the asymptotic value $0.75$ reached in the limit of infinite temperature owing to the threefold degeneracy of the triplet state in zero magnetic field.
A comparison of the dimer-triplet density obtained for $N_{s}=30$ and $N_{s}=40$ spins from exact diagonalization (ED) reveals small finite-size effects. ED for $N_{s}=40$ spins provides an interpolation between the spin-star decoupling approximation and the EMD model at low temperatures and QMC at high temperatures. The excitation from a dimer-singlet state to a dimer-triplet state is neglected within the EMD model and the dimer-triplet density should accordingly always decrease with increasing temperature. On the other hand, the local minimum observed in the dimer-triplet density arises from the competition of the lower excitation energy for the dimer-singlet state and the higher degeneracy of the triplet state. In fact, the lowest states not taken into account by the EMD model are at energies slightly above $0.6J_{1}$ for $J_{2}=1.7J_{1}$ and these are responsible for the triplet density increasing again with increasing temperature after going through a minimum. However, the spin-star decoupling approximation takes only some of these into account [as reflected also by the lower entropy $s$ compared to ED and QMC, see Fig. 8(b)] and approximates their energies, thus explaining the deviation from $N_{s}=40$ ED and QMC, which are in good agreement for $T/J_{1}\gtrsim 0.2$.
@@
Figure 8(c) displays the specific heat per spin $c$ as a function of temperature for the same value of the interaction ratio. A notable feature is the presence of a low-temperature shoulder, superimposed on the high-temperature maximum of the specific heat, as seen clearly in ED and QMC results. As anticipated, the spin-star and EMD approaches fail to capture the behavior of the specific heat at higher temperatures as they neglect high-energy excitations above the DT phase.
On the other hand, these approaches accurately account for low-lying excitations which leads to a close agreement between the ED data for $N_{s}=40$ and the analytical results obtained from the spin-star decoupling approximation at low enough temperatures. A comparison of the peak height in the specific heat obtained from ED and QMC reveals relatively large finite-size effects at low temperatures. The EMD model provides further insight into the behavior of the specific heat at low temperatures, which effectively bridges the gap between the available QMC data ($T/J_{1}>0.1$) and extends it further towards lower temperatures. Figure 8(c) shows that thermal activation of the specific heat indeed occurs at progressively lower temperatures as the system size increases such that the EMD model provides the best approximation to the thermodynamic limit at low temperatures.
@@
We now turn to a discussion of the dimer-triplet density, entropy, and specific heat in zero magnetic field for the case $J_{2}/J_{1}=1.5$ plotted in Fig. 9. Here the numerical data for the dimer-triplet density exhibits only a shallow minimum, which is consistent with the gap for one triplet less being only slightly smaller than that for an additional triplet at $J_{2}/J_{1}=1.5$, compare Fig. 7. By contrast, the spin-star decoupling approximation fails to account for this minimum. The reason is that, within the spin-star decoupling approximation, the excitation energies from the DT phase to singlets and triplets are equal for this interaction ratio. Consequently, the spin-star approach cannot explain the thermal population into singlet states for the interaction ratio $J_{2}/J_{1}\leq 1.5$. Comparing the results obtained from QMC simulations for a system size of $6\times 6$ (total of 180 spins) with those from ED for 40 spins reveals nearly perfect agreement, indicating negligible finite-size effects. However, the discrepancy between ED results for 30 spins and 40 spins at intermediate temperatures ($T/J_{1}>0.1$) may stem from artificial symmetries imposed for the system size of 30 spins by considering periodic boundary conditions, which are absent for the system size of 40 spins.
The entropy and specific heat for the interaction ratio $J_{2}/J_{1}=1.5$ in zero magnetic field are shown in Figs. 9(b) and 9(c), respectively, and exhibit qualitatively similar behavior to that for the interaction ratio $J_{2}/J_{1}=1.7$. In the high-temperature regime, small finite-size effects are observed. While the high-temperature peak of the specific heat shifts slightly towards higher values with increasing system size, the low-temperature peak is conversely suppressed towards lower values. The results obtained from the spin-star decoupling approximation for 40 spins are slightly overestimated compared to ED data for $N_{s}=40$ at low temperatures $T/J_{1}\lesssim 0.1$. The discrepancy at higher temperatures arises because of the presence of low-energy excitations to states that share the same energy as the Lieb-Mattis ferrimagnetic state at zero magnetic field (compare the discussion in Sec. III.3), which is not accurately accounted for by the spin-star decoupling approximation simply ignoring the noncommutativity of the spin-star Hamiltonians. The QMC data are consistent with the intriguing double-peak temperature dependence of the zero-field specific heat with low- and high-temperature peaks located around $T/J_{1}\approx 0.12$ and 0.6, respectively.
@@
With a yet lower value of the interaction ratio $J_{2}/J_{1}$, the analytical approaches provided by the spin-star and EMD models neglect more and more important low-energy excitations. For the interaction ratio $J_{2}/J_{1}$=1.3, both of these approaches only provide an accurate description at very low temperatures as evidenced by the temperature dependencies of the dimer-triplet density, entropy, and specific heat depicted in Fig. 10. This limitation arises from the failure of these analytical approaches to accurately capture the degeneracy of the LM phase in zero magnetic field with possible values of the $z$ component of the total spin $S_{T}^{z}=-S_{T},-S_{T}+1,...,S_{T}$, and the low-energy excitations to these states are not properly accounted for either. The dimer-triplet density monotonically increases without any local minimum for the interaction ratio $J_{2}/J_{1}$=1.3, as evidenced by the ED data for $N_{s}=40$ spins. This behavior is caused by the lowest excitation in this parameter regime arising from an additional dimer-triplet, compare Fig. 7. The EMD approach neglects the lowest excitations and thus only gets the $T=0$ limit right, but fails to capture the leading low-temperature asymptotics. By contrast, the star approach does account for these excitations, even if only in an approximate fashion, and thus captures at least qualitatively the low-temperature behavior of the triplet density.
@@ IV.2 Magnetization, entropy, and specific heat under magnetic field
It can be observed from Fig. 11(b) that the entropy starts from its nonzero residual value, which is progressively converging with increasing system size to the residual entropy of the dimer model on the square lattice per unit cell $s_{d}\approx 0.29156$ (per spin $\approx 0.058312$) [64, 65, 43]. Contrary to this, one detects a sizable entropy gain reaching the specific value $s=0.13256$ at the transition field between the DT and MD phases, which is almost independent of the system size as corroborated by ED, the EMD model, as well as the spin-star decoupling approximation. Similar qualitative agreement between results of all three aforementioned methods is observed in the peculiar asymmetric double-peak structure of the specific heat, which can be detected in the vicinity of the magnetic-field-driven phase transition between DT and MD phases [see the inset of Fig. 11(c)]. The lower peak height observed below the relevant transition field can be attributed to the macroscopic degeneracy of the DT ground state, which is not lifted by the magnetic field in contrast to the MD phase that becomes non-degenerate at any nonzero magnetic field.
@@
The results obtained from the spin-star decoupling approximation are in very good agreement with the ED data even for magnetic fields higher than $h/J_{1}\gtrsim 1.0$, up to the end of the Lieb-Mattis ferrimagnetic phase, and even at moderately high temperatures $T/J_{1}=0.1$ still constitute a good approximation of the field dependence of the entropy and specific heat, as depicted in Figs. 11(b) and 11(c).
@@
Figure 12: (a) Magnetization; (b) entropy; (c) specific heat for the interaction ratio $J_{2}/J_{1}=1.5$ as a function of magnetic field at two different values of temperature $T/J_{1}=0.01$ and $T/J_{1}=0.1$. The inset in (a) shows $T=0$ ED and DMRG results.
@@
Figure 12(a) shows the magnetization curves for the interaction ratio $J_{2}/J_{1}=1.5$. There is a distinct magnetization jump from the DT phase to the MD phase at zero temperature, as shown in the inset of Fig. 12(a). However, the 1/5-plateau is sufficiently small to be smeared out at finite temperatures. Consequently, there is a sharp rise in magnetization from the 0-plateau to the 3/5-plateau even at temperatures as low as $T/J_{1}=0.01$, which becomes smoother with increasing temperature. Owing to the presence of the MD phase within a narrow range of magnetic field strengths, the EMD model is in excellent agreement with the ED data up to a magnetic field of approximately $h/J_{1}\approx 0.54$ at a very low temperature $T/J_{1}=0.01$, although it fails to account for further increases in magnetization with increasing magnetic-field strength. On the other hand, the spin-star decoupling approximation is still capable of qualitatively reproducing the ED data even at higher magnetic fields. The comparison of ED data for $N_{s}=40$ spins with QMC data for 36 unit cells (i.e., $N_{s}=180$ spins) indicates negligible finite-size effect in this parameter region.
@@
The results for the low-temperature entropy exhibit a relevant discrepancy between the ED and spin-star decoupling approximation only in the close vicinity of the transition field $h/J_{1}=0.5$ [see Fig. 12(b)]. The excessive macroscopic degeneracy predicted by the spin-star decoupling approximation relates to a triple phase coexistence point of the LM, DT, and MD phases, which is located exactly at the magnetic field $h/J_{1}=0.5$ for $J_{2}/J_{1}=1.5$. Recall that the triple point obtained within the spin-star approach is slightly shifted towards a lower value of the magnetic field with respect to its true coordinates [see Fig. 2(b)]. At a higher temperature $T/J_{1}=0.1$, the discrepancy between ED and the spin-star decoupling approximation is significantly reduced, because all relevant low-lying excited states are thermally activated. Contrary to this, we see the opposite effect for the EMD model with much lower entropy, because high-energy excitations are missing within this description. A similar effect can also be observed in the field dependence of the specific heat [Fig. 12(c)]. While the EMD model reproduces the ED data at low temperatures, the spin-star decoupling approximation correctly describes the double peak of the specific heat also at a higher temperature of $T/J_{1}=0.1$. At this higher temperature, the double-peak structure of the specific heat is also confirmed by the QMC simulations for the bigger system of $6\times 6$ unit cells.
@@
Last but not least, magnetization, entropy, and specific heat as a function of temperature are depicted in Fig. 13 for the interaction ratio $J_{2}/J_{1}=1.3$. As discussed in Sec. III.3, for the interaction ratio $J_{2}/J_{1}\lesssim 1.4$, the lowest excitation above the DT phase corresponds to the creation of an additional triplet [see Fig. 7]. For this reason, the EMD model is unable to describe the finite-temperature properties for the interaction ratio $J_{2}/J_{1}=1.3$. On the other hand, our second analytical approach derived from the spin-star decoupling approximation provides a quite reliable description of all quantities above the transition field $h/J_{1}=0.3$. The most important artifacts of the spin-star decoupling approximation are just below the transition field and at $T/J_{1}=0.01$ where it yields a spike in the entropy $s$ [Fig. 13(b)] and a second maximum in the specific heat $c$ [Fig. 13(c)] while such pronounced maxima are absent in the ED results. The reason for this difference is that the field-induced DT-LM transition is a true first-order transition in the full model such that excitations are scattered over a certain energy range below the transition while the spin-star decoupling approximation treats these as a degenerate manifold that comes down at the transition field. Conversely, a temperature $T/J_{1}=0.1$ is sufficiently high to thermally excite these low-lying excitations in the full model such that at this higher temperature the $N_{s}=40$ ED results also exhibit a pronounced maximum below the transition field, both in the entropy [Fig. 13(b)] and the specific heat [Fig. 13(c)]. The latter behavior is confirmed by our QMC simulations on the bigger system of $6\times 6$ unit cells.